randy has a jar containing 51 coins, all quarters and nickels. totaling $5.55. how many quarters are in the jar?
Answer by richwmiller(17219)

q+n=51
n=51-q
25q+5n=555
5q+n=111
5q+51-q=111
4q=60
q=15
n=36